# Pre-training Differentially Private Models with Limited Public Data

### v1

@@ 3.3 DP fine-tuning is robust to noise
Empirical evidence has shown that DP fine-tuning is comparable to (though worse than) the standard fine-tuning [88, 48, 23, 57, 9, 11, 12], despite that $\sigma\neq 0$. Such a phenomenon implies that comparing public and DP finetuning, we have $\Delta^{\star}_{\textup{pub}}(B)\approx\Delta^{\star}_{\textup{priv}}(B)$; That is, the decelerator becomes small after the public pre-training. This is conceptually illustrated in Figure 5 (right column, say $B=100$), where the DP curve is close to the non-DP curve at moderate $B$ during fine-tuning, but not so during pre-training.
To understand the stark contrast between DP fine-tuning and DP pre-training, we plug in the optimal $B^{\star}_{\text{DP}}=\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$ from 2.4 to $\Delta^{\star}_{\textup{priv}}(B)$. Then, we have the optimal improvement of DP-SGD as
$\Delta^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}})=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{2\sqrt{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\cdot\sigma^{2}\textup{tr}({\mathbf{H}})/c^{2}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}.$
Notice that by choosing $B_{\text{non-DP}}=2B^{\star}_{\text{DP}}$, we have
$\Delta^{\star}_{\textup{pub}}(2B^{\star}_{\text{DP}})=\Delta^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}}).$ (11)
This indicates that DP-SGD with the optimal batch size is as fast as SGD with twice the batch size. Moreover, if $B^{\star}_{\text{DP}}$ is moderate, then DP-SGD can converge at similar speed to the SGD with optimal batch size.
@@
From (9), we observe that non-DP training is data-efficient only if $B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\ll\textup{tr}({\mathbf{H}}\mathbf{\Sigma})$, otherwise, we can decrease $B_{\text{non-DP}}$ to effectively improve $\Delta^{\star}_{\textup{pub}}$. Therefore, by (11), DP training is data-efficient only if $B^{\star}_{\text{DP}}=\frac{1}{2}B_{\text{non-DP}}\ll\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{2{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}$, which holds in fine-tuning but not in early stage of pre-training (see Figure 6 right column, which illustrates the magnitude of the above three terms for pre-training and fine-tuning stages).
Figure 6: Evolution of terms in (7) and (9) that explains the deceleration of DP optimization, during pre-training (upper) and fine-tuning (lower) ViT-Base on CIFAR100.
@@ 3.4 Extension to general optimizers
The analysis in the previous two sections can be easily extended to arbitrary optimizers as well as techniques such as weight decay, gradient clipping, and parameter-efficient fine-tuning (PEFT). Let $p$ be an optimizer’s post-processor of gradient and consider
${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta p({\mathbf{g}}_{t}).$
@@
$\text{SGD$(\mu,\lambda)$:}\quad p({\mathbf{g}};{\mathbf{b}},{\mathbf{w}})=\mu{\mathbf{b}}+{\mathbf{g}}+\lambda{\mathbf{w}}.$
Similarly, we can write any PEFT for any optimizer, e.g.
@@
where ${\mathbf{M}}\in\{0,1\}$ is an element-wise mask, with ${\mathbf{M}}=0$ indicating that the parameter is non-trainable or frozen.
With respect to different $p(\cdot)$, the expected per-iteration loss improvement becomes
$\Delta L(\eta)$ $=\eta{\mathbf{G}}^{\top}\mathbb{E}[p({\mathbf{g}})]-\frac{\eta^{2}}{2}\mathbb{E}[p({\mathbf{g}})^{\top}{\mathbf{H}}p({\mathbf{g}})]).$ (12)
Applying Assumption 2.1 and the optimal learning rate, we can derive the expected per-sample per-iteration improvement, which is left in Section A.5 due to the complication in its form.
Specially, we consider optimizers such that $p(\cdot)$ is scale-invariant (i.e. $p(c{\mathbf{G}})=p({\mathbf{G}})$), such as SignSGD/Adam [4] or normalized SGD/LAMB [60, 54, 86].
Implication 3.2.
@@ 4.1 Necessity of public data in DP pre-training
In this section, we propose the DP continual pre-training strategy and demonstrate that the deceleration by DP can be effectively mitigated by using a certain amount of public data. We consider the mixed data training that uses both public data (with subscript ${}_{0}$ for related hyperparameters) and private data (with subscript ${}_{1}$). Then (4) becomes
@@
From the optimal choice, we see $\alpha^{\star}\neq 0$. It indicates that with the ability to access any public data (${\mathbf{g}}_{j,t}$), the optimal improvement of mixed data training is larger than private training, i.e., $\Delta L^{\star}_{\text{OnlyPrivate}}<\Delta L^{\star}_{\textup{mixed}}.$
@@
Similarly, more private data also improves upon public training. The optimal $\alpha^{\star}\neq 1$ also indicates that $\Delta L^{\star}_{\text{OnlyPublic}}<\Delta L^{\star}_{\textup{mixed}}$ i.e., the improvement of mixed data training is larger than public training.
4.2 DP Continual pre-training strategy overview
Motivated by Section 3 and Remark 4.1, we propose a two-stage DP continual pre-training strategy: public pre-training for $sT$ steps followed by private continual pre-training for $(1-s)T$ steps, which is equivalent to setting $\alpha_{t}=\mathbb{I}(t<sT)$ and the constant $10<s<$ 1 controls the portion of steps of public training.
Remark 4.3.
Although (13) suggests that $\alpha_{t}\in(0,1)$, we only set binary $\alpha_{t}\in\{0,1\}$ to avoid the difficulty of implementing two data loaders and two back-propagation mechanisms simultaneously (DP v.s. standard) in large optimization, e.g. memory and synchronization issues.
We highlight that DP continual pre-training is as distinct from DP fine-tuning as their non-DP counter-parts, despite both methods extending the learning from a pre-trained stage: the continual pre-training learns common knowledge without adapting to a specific task and serves a richer foundation for many downstream tasks (see more discussion in Appendix C). In fact, we show in Section A.8 that the loss improvement of DP continual pre-training can be almost as fast as non-DP pre-training (FullyPublic), thus closing the utility gap.
Table 3: Optimization by different training strategies.
public data private data privacy
OnlyPublic SGD not used Yes
OnlyPrivate not used DP-SGD Yes
FullyPublic SGD SGD No
FullyPrivate DP-SGD DP-SGD Yes
Mixed (${\mathbf{g}}_{\alpha}$) SGD DP-SGD Yes
4.3 Technical Details: Stage Switching
We now demonstrate the technical details of our approach. When to switch: The switching, i.e. selecting $s$, can be automatically determined by public statistics without using any DP budget. For example, we can use early stopping based on the loss, i.e., switch after the loss on the validation dataset stops decreasing (see Figure 7, SwitchPoint of two Mixed training at different steps) or accuracy to switch to DP training. Alternatively, we can monitor $B^{\star}_{\text{DP}}=\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$ and switch when it drops to the same level of $B_{\text{non-DP}}$, so that $\Delta^{\star}_{\textup{priv}}\approx\Delta^{\star}_{\textup{pub}}$.
Figure 7: Convergence of GPT2-small on CodeParrot with different pre-training strategies ($\epsilon=8$ if Mixed).
How to switch Depending on whether the optimizer is scale-invariant, e.g. SGD v.s. Adam, the learning rate may need rescaling after switching to DP optimizers, in order to avoid training instability or even failure.
Suppose that on (10) is small, we recognize that DP-SGD needs to rescale the learning rate of SGD by $1/c$, where $c$ is per-sample clipping factor. However, DP-Adam(SignSGD) and DP-LAMB do not need to. The proof of these facts is provided in Section A.6.
@@ 5 DP vision foundation models on ImageNet
We leverage the DP continual pre-training strategy discussed in Section 4 to train vision transformers [26] on ImageNet datasets. We use ImageNet-1k (1.3M images, 1k classes; [24]) for public pre-training, then ImageNet-11k (formally known as ImageNet-21k-P [^3] ImageNet-11k improves the dataset quality and thus the resulting models (see Table 3 in [69]), by removing the infrequent classes from the full ImageNet-21k, thus remaining 78% of its 14M images and 48% of 21.8k classes., 11M images, 11k classes; [69]) for private pre-training. Notice that ImageNet-11k/21k is significantly harder to learn, with SOTA accuracy $\approx 47\%$ [69, 73] as compared to $>85\%$ for ImageNet-1k. We apply data augmentation including random flip, contract shift, and rotation, and resizing to $224\times 224$. We evaluate our DP ViT on upstream and downstream tasks (including few-shot), achieving high accuracy under low computation budget (compared to existing DP pre-trained models in Table 4) while protecting the privacy of 90% training data.
Table 4: Pre-training strategies of models. Standard non-DP training is marked in black; DP training is in green. $\dagger$ indicates self-supervised without using the labels. images $\times$ is the total number of images used (dataset size×$\times$×epochs).
@@
(Syn)ViP [89] ViT-Base ${}^{\dagger}$ Shaders21k 1.3B
ViP [89] ViT-Base ${}^{\dagger}$ ${}^{\dagger}$ Shaders21k 1.9B
DINO [18] ViT-Base ${}^{\dagger}$ ImageNet-1k 0.3B
Ours This work ViT-Base ${}^{\dagger}$ ImageNet-11k ImageNet-1k 0.7B
@@ 5.1 Training strategy
For the public pre-training, we follow the self-supervised learning by [18] (self-distillation with no labels, or DINO), whereas the private continual pre-training is a supervised learning following [69] [^4] We observe that, using augreg (a supervised pre-training strategy) for public pre-training on ImageNet-1k and/or ImageNet-21k for private pre-training gives very similar result.. We employ AdamW optimizer with batch size $B=4096$ and learning rate $\eta=0.0002$ set by the line search [^5] Our training strategy is similar to a concurrent work [89], with critical differences highlighted in Appendix C..
When automatically switching with early stopping (c.f. Section 4.3) from public to private pre-training, the classification head (the last layer) is re-initialized, after which we train with DP linear probing for 10 epochs then with DP full-parameter training for 20 epochs, with a total $\approx$100k training steps. This strategy achieves an upstream accuracy 41.5% on ImageNet-11k by our ViT-Base under $\epsilon=8$, and 39.8% under $\epsilon=2$.
@@ 5.2 Downstream performance
[…]
In Table 7, we extend the evaluation of full fine-tuning and linear probing to SOTA non-DP baselines and to million-image scale Our DP model achieves 55.7% (+9.6% over ViP) on Places365 with 1.8M images and 60.0% (+21.9% over ViP) on iNat2021 with 2.7M images. The current non-DP SOTA is 57-60% on Places365 [23, 26] and about 64% on iNat2021 [78, 59] after pre-training on $2.7\sim 4\textup{B}$ images. This showcases the effectiveness of DP pre-training (only leveraging 0.7B images).
Table 7: Linear-probing accuracy (non-DP) of pre-trained models, except “full” indicating full-parameter. For ImageNet-1k (seen in our pre-training), it is only meaningful to compare our models to DINO and observe the benefit of continual pre-training, but not to others.
@@
[…]
@@ 6 Discussion
In this paper, we conduct an insightful and unified convergence analysis on DP optimization. Specifically, we identify the decelerator (10) of DP training as a result of the per-sample gradient clipping, the noise and the Hessian, which can be significantly mitigated by a small amount ($<10\%$) of public training. Consequently, we propose DP continual pre-training to enrich a foundation model’s knowledge via the private data. We are excited to demonstrate that, on large-scale systems, our DP pre-training is almost as accurate and implementable as the fully public pre-training.

### v2

@@ 3.3 DP fine-tuning is robust to noise
Empirical evidence has shown that DP fine-tuning is comparable to (though slightly worse than) the standard non-DP fine-tuning [89, 49, 24, 58, 10, 14, 12], despite that $\sigma\neq 0$. Such a phenomenon implies that comparing public and DP finetuning, we have $\Delta L^{\star}_{\textup{pub}}(B)\approx\Delta L^{\star}_{\textup{priv}}(B)$. That is, the decelerator becomes small after the public pre-training. This is conceptually illustrated in Figure 5(b), where the DP curve is close to the non-DP curve at moderate $B$ during fine-tuning, but not so during pre-training.
To understand the stark contrast between DP fine-tuning and DP pre-training, we plug in the optimal $B^{\star}_{\text{DP}}=\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$ from 2.4 to $\Delta L^{\star}_{\textup{priv}}(B)$. Then, we have the optimal improvement of DP-SGD as
$\Delta L^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}})=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{2\sqrt{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\cdot\sigma^{2}\textup{tr}({\mathbf{H}})/c^{2}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}.$
Implication 3.1.
Notice that by choosing $B_{\text{non-DP}}=2B^{\star}_{\text{DP}}$ in (7), we have that DP-SGD with the optimal batch size is as fast as the standard SGD with twice the batch size,
$\Delta L^{\star}_{\textup{pub}}(2B^{\star}_{\text{DP}})=\Delta L^{\star}_{\textup{priv}}(B^{\star}_{\text{DP}}).$ (9)
Moreover, if $B^{\star}_{\text{DP}}$ is moderate, then DP-SGD can converge at similar speed to a fast converging SGD that uses a moderate batch size.
@@
From (7), we observe that non-DP training is data-efficient only if $B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\ll\textup{tr}({\mathbf{H}}\mathbf{\Sigma})$. Otherwise, we can decrease $B_{\text{non-DP}}$ to effectively improve $\Delta L^{\star}_{\textup{pub}}$. Therefore, DP training is data-efficient only if $B^{\star}_{\text{DP}}\approx\frac{1}{2}B_{\text{non-DP}}\ll\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{2{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}$, which holds in fine-tuning but not in early stage of pre-training [^3] An intuitive explanation is that $\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}=\frac{\mathbb{E}[({\mathbf{g}}_{i}-{\mathbf{G}})^{\top}{\mathbf{H}}({\mathbf{g}}_{i}-{\mathbf{G}})]}{{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}=\frac{\mathbb{E}({\mathbf{g}}_{i}^{\top}{\mathbf{H}}{\mathbf{g}}_{i})}{\mathbb{E}({\mathbf{g}}_{i})^{\top}{\mathbf{H}}\mathbb{E}({\mathbf{g}}_{i})}-1$ resembles the variance of per-sample ${\mathbf{g}}_{i}$ in the space of ${\mathbf{H}}$, which decreases as the model learns the common representation. We illustrate the magnitude of the three terms in (6) for pre-training and fine-tuning stages in Figure 6.
Figure 6: Evolution of terms in (6) and (7) that explains the deceleration of DP optimization, during pre-training (left two) and fine-tuning (right two) ViT-Base on CIFAR100.
In the fine-tuning phase of Figure 6, $\textup{tr}({\mathbf{H}})$ quickly decreases and so does the decelerator. Hence a moderate $B_{\text{DP}}\approx 100$ can allow fast convergence. However, in the pre-training phase, $\textup{tr}({\mathbf{H}})$ increases to a large value within 5 epochs and remains for a long time (say epoch 5 to 40) before it decreases again. Consequently, DP convergence is initially fast but only for a short period and overall DP optimization is much slower than non-DP optimization, as shown in Figure 1(a)(c).
@@ 3.4 Extension to general optimizers
The analysis in the previous two sections can be easily extended to arbitrary optimizers as well as techniques such as weight decay, gradient clipping, and parameter-efficient fine-tuning (PEFT). Let $p$ be an optimizer’s post-processor of gradient and consider ${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta p({\mathbf{g}}_{t}).$
@@
Specially, we consider optimizers such that $p(\cdot)$ is scale-invariant (i.e. $p(c{\mathbf{G}})=p({\mathbf{G}})$), such as SignSGD/Adam [4] or normalized SGD/LAMB [61, 55, 87]. Applying Assumption 2.1 and the optimal learning rate, we derive the expected per-sample per-iteration improvement in 3.3, leaving the details Section A.5.
Implication 3.3.
@@ 4.1 Necessity of public data in DP pre-training
In this section, we propose the DP continual pre-training strategy and demonstrate that the deceleration by DP can be effectively mitigated by using a certain amount of public data. We consider the mixed data training that uses both public data (with subscript 0 for related hyperparameters) and private data (with subscript 1). Then SGD becomes
@@
We see that the optimal choice in (10) gives $\alpha^{\star}\neq 0$, which indicates that the optimal improvement of mixed data training is larger than private training, i.e., $\Delta L^{\star}_{\text{OnlyPrivate}}<\Delta L^{\star}_{\textup{mixed}}$ and the public data helps. On the other hand, the fact that optimal $\alpha^{\star}\neq 1$ also indicates that $\Delta L^{\star}_{\text{OnlyPublic}}<\Delta L^{\star}_{\textup{mixed}}$ i.e., the private data helps.
4.2 DP Continual pre-training strategy overview
Motivated by Section 3 and Remark 4.1, we propose a two-phase DP continual pre-training strategy in Algorithm 1: public pre-training for $sT$ steps followed by private continual pre-training for $(1-s)T$ steps, which is equivalent to setting $\alpha_{t}=\mathbb{I}(t<sT)$ and the constant $10<s<$ 1 controls the portion of steps of public training.
@@ 4.2 DP Continual pre-training strategy overview
Although (10) suggests that $\alpha_{t}\in(0,1)$, we only set binary $\alpha_{t}\in\{0,1\}$ so as to not implement two data loaders and two back-propagation mechanisms simultaneously (DP v.s. standard). Note the methods in Table 2 are difficult to scale and not yet openly implemented on distributed systems due to memory and synchronization issues.
Figure 7: Convergence of GPT2-small on CodeParrot with different pre-training strategies ($\epsilon=8$).
We highlight that DP continual pre-training is as distinct from DP fine-tuning as their non-DP counter-parts, despite both methods extending the learning from a pre-trained phase: the continual pre-training learns common knowledge without adapting to a specific task and serves a richer foundation for many downstream tasks (see more discussion in Appendix C). In fact, we show in Section A.7 that the loss improvement of DP continual pre-training can be almost as fast as non-DP pre-training (FullyPublic), thus closing the utility gap.
The switching, i.e. selecting $s$, can be automatically determined by public statistics without using any DP budget. For example, we can use early stopping based on the loss or accuracy, i.e., we switch after the metrics on the validation dataset stop improving (see Figure 7). Alternatively, we can monitor $B^{\star}_{\text{DP}}$ and switch when it drops to a moderate value so that the DP training converges fast by (9).
@@ 5 DP vision foundation models on ImageNet
We leverage the DP continual pre-training strategy discussed in Section 4 to train vision transformers [27] on ImageNet datasets. We use ImageNet-1k (1.3M images, 1k classes; [25]) for public pre-training, then ImageNet-11k (formally known as ImageNet-21k-P [^4] ImageNet-11k improves the dataset quality and thus the resulting models (see Table 3 in [70]), by removing the infrequent classes from the full ImageNet-21k, thus remaining 78% of its 14M images and 48% of 21.8k classes., 11M images, 11k classes; [70]) for private pre-training. Notice that ImageNet-11k/21k is significantly harder to learn, with SOTA accuracy $\approx 47\%$ [70, 74] as compared to $>85\%$ for ImageNet-1k. We apply data augmentation including random flipping, contract shift, rotation, and resizing to $224\times 224$.
We evaluate our DP ViT on upstream and downstream tasks (including few-shot), achieving high accuracy under low computation budget (compared to existing DP pre-trained models in Table 3). In short, we show that unlocking more data (90% of full ImageNet), which cannot be used by DINO due to privacy concern, is significantly beneficial.
Table 3: Pre-training strategies of models. Standard non-DP training is marked in black; DP training is in green. $\dagger$ indicates self-supervised without using the labels. “Images $\times$” is the total number of images used (dataset size×$\times$×epochs). “Non-privacy” means no DP guarantee on a subset of training data due to the non-DP pre-training phase.
@@
(Syn)ViP [90] ViT-Base Shaders21k Shaders21k 1.3B
ViP [90] ViT-Base Shaders21k LAION400M Shaders21k 1.9B
DINO [19] ViT-Base ImageNet-1k ImageNet-1k 0.3B
Ours This work ViT-Base ImageNet-1k ImageNet-11k ImageNet-1k 0.7B
@@ 5.1 Training strategy
For the public pre-training, we follow the self-supervised learning by [19] (self-distillation with no labels, or DINO), whereas the private continual pre-training is a supervised learning following [70] [^5] We observe that, using augreg (a supervised pre-training strategy) for public pre-training on ImageNet-1k and/or ImageNet-21k for private pre-training gives very similar result.. We employ AdamW optimizer with batch size $B=4096$ and learning rate $\eta=0.0002$ set by the line search. Our training strategy is similar to a concurrent work [90], with critical differences highlighted in Appendix C.
When automatically switching from public to private pre-training, the classification head (the last layer) is re-initialized because the number of classes is different in the two pre-training phases. This switching is triggered by early stopping. In the continual pre-training phase, we train with DP linear probing for 10 epochs then with DP full-parameter training for 20 epochs, with a total $\approx$100k training steps. This strategy achieves an upstream accuracy 41.5% on ImageNet-11k by our ViT-Base under $\epsilon=8$, and 39.8% under $\epsilon=2$.
5.2 Algorithm implementation
We employ fastDP library to apply the DP optimizers with automatic per-sample clipping [14] and layer-wise clipping style [9]. Specifically, the DP optimization is under the multi-GPU distributed system, using DP-ZeRO [7] and mixed-precision training. We calibrate the DP noise using the improved Renyi accountant.
@@ 5.3 Downstream performance
Oursϵ=2 36.42 48.27 64.74 74.62 100
Oursϵ=8 42.57 57.15 65.26 76.38 100
Table 6: Linear-probing accuracy (non-DP) of pre-trained models, except “full” indicating full-parameter.
@@
Oursϵ=2 76.2 52.5 46.5 8 / 5 / 10
Oursϵ=8 77.9 53.0 49.1 8 / 5 / 10
Oursϵ=2(full) 78.0 55.6 57.2 8 / 5 / 10
Oursϵ=8(full) 78.5 55.7 60.0 8 / 5 / 10
@@
In Table $6$For ImageNet-1k, we strike through some results because this dataset was used in the pre-training. It is only meaningful to compare our models to DINO and observe the benefit of continual pre-training, but not to others., we extend the evaluation of full fine-tuning and linear probing to SOTA non-DP baselines and to million-image scale Our DP model achieves 55.7% (+9.6% over ViP) on Places365 with 1.8M images and 60.0% (+21.9% over ViP) on iNat2021 with 2.7M images. The current non-DP SOTA is 57-60% on Places365 [24, 27] and about 64% on iNat2021 [80, 60] after pre-training on $2.7\sim 4\textup{B}$ images. This showcases the effectiveness of DP pre-training as our models only leverage 0.7B images.
@@ 6 Discussion
In this paper, we conduct an insightful and unified convergence analysis on DP optimization. Specifically, we identify the decelerator (8) of DP training as a result of the per-sample gradient clipping, the noise and the Hessian, which can be significantly mitigated by a small amount ($<10\%$) of public training. Consequently, we propose DP continual pre-training that is almost as accurate and implementable as the fully public pre-training.
Acknowledgement
The work of Xinwei Zhang was partially done while interning at Amazon. Mingyi Hong holds concurrent appointments as an Amazon Scholar and as a faculty at the University of Minnesota. This paper describes their work performed at Amazon.